Using High-Low to Calculate Fixed Cost, Calculate the Variable Rate, and Construct a Cost Function

Pizza Vesuvio makes specialty pizzas. Data for the past 8 months were collected:

Month Labor Cost($) Employee Hours
January 9,490 460
February 7,000 530
March 7,531 560
April 8,140 420
May 9,787 490
June 8,490 390
July 11,868 630
August 7,500 350
Pizza Vesuvio's controller wants to calculate the fixed and variable costs associated with labor used in the restaurant.

In your calculations, round the variable rate per employee hour to the nearest cent. If required, round your final answers to the nearest cent.

Required:

1. Using the high-low method, calculate the fixed cost of labor.
$
915
2. Using the high-low method, calculate the variable rate.
$
17.39
per employee hour

3. Using the high-low method, construct the cost formula for total labor cost.

Total labor cost = $
915
+ ($
17.39
× Employee hours)
1
Expert's answer
2020-09-21T16:29:31-0400

2. "Variable\ Rate\ Per\ Employee="


"=\dfrac{High\ Cost-Low\ Cost}{High\ Employee\ Hours-Low\ Employee\ Hours}="


"=\dfrac{\$11868-\$7500}{630-350}=\$15.60\ per \ employee\ hour"



1. "Fixed\ Cost="


"=Total\ Cost-(Variable\ Rate\times Employee\ Hours)="


"=\$ 11868-(\$ 15.60\times630)="


"=\$ 7500-(\$ 15.60\times350)=\$ 2040"


3. "Total\ Labor\ Cost="


"=\$ 2040+(\$ 15.60\times Employee\ Hours)"